40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 19 23:44:00 2018, gautam, Update, IOO, IMC loop checkup IMC_RF_noise.pdfIMC_OLTF_20180320.pdf
    Reply  Tue Mar 20 16:53:03 2018, gautam, Update, IOO, IMC loop checkup IMC_RF_noise.pdf
       Reply  Tue Mar 20 21:08:03 2018, gautam, Update, IOO, IMC loop checkup 
          Reply  Tue Mar 20 22:44:45 2018, gautam, Update, IOO, IMC loop checkup 
             Reply  Wed Mar 21 15:52:45 2018, gautam, Update, IOO, MC error point calibration IMC_PDHdisc_20180321.pdfIMC_PDHdisc785_20180321.pdfIMC_oscope.pdf
                Reply  Wed Mar 21 17:31:21 2018, gautam, Update, IOO, MC error point calibration IMC_RF_noise_calib.pdf
                   Reply  Wed Mar 21 21:13:44 2018, gautam, Update, IOO, IMC noise budget IMC_RF_noise_calib.pdf
                      Reply  Tue Mar 27 01:39:44 2018, gautam, Update, IOO, PSL noise eater was off IR_ALS_20180326.pdf
                         Reply  Tue Mar 27 19:32:03 2018, arijit, Update, IOO, PSL noise eater was off MCREFL_TF.pdfMCREFL_SPECTRUM.pdf
                            Reply  Tue Mar 27 23:37:35 2018, gautam, Update, IOO, MC REFL PD removed IMG_6955.JPG
                               Reply  Wed Mar 28 21:31:39 2018, gautam, Update, IOO, MC REFL PD removed 
                            Reply  Wed Mar 28 21:47:37 2018, arijit, Update, IOO, MCREFL_PD Optical response measurement LISO_fit_with_info.pdfLISO_fit.zip
                               Reply  Thu Mar 29 17:57:36 2018, arijit, Update, IOO, MCREFL_PD Optical response measurement NOISE_EATER_On_OFF.pdf
                                  Reply  Fri Mar 30 06:14:31 2018, rana, Update, IOO, MCREFL_PD Optical response measurement 
                                     Reply  Fri Mar 30 22:37:36 2018, Kevin, Update, IOO, MCREFL_PD Optical response measurement REFLPD_DemodBoard.pdf
                      Reply  Thu Apr 5 04:36:56 2018, Kevin, Update, IOO, Coil driver noise MC3.pdfMC2.pdfCoilDriver.pdf
                         Reply  Thu Apr 5 12:13:18 2018, Koji, Update, IOO, Coil driver noise 
                            Reply  Fri Apr 6 21:39:09 2018, gautam, Update, IOO, MC2 suspension health checkup MC2_Freeswinging.pdf
                               Reply  Mon Apr 9 18:46:03 2018, gautam, Update, IOO, Further debugging MC_Freeswinging.pdfIMC_NB_20180409.pdf
                                  Reply  Tue Apr 10 14:28:44 2018, gautam, Update, IOO, Further debugging 
                                     Reply  Wed Apr 11 01:34:31 2018, gautam, Update, IOO, Activities today 
                                        Reply  Thu Apr 12 18:12:49 2018, gautam, Update, ALS, NPRO channels hijacked IR_ALS_20180412.pdfBeatMouthDrift.pngETMX_20180416.jpg
                                           Reply  Fri Apr 13 11:02:41 2018, gautam, Update, ALS, EX fiber polarization drift EX_pol_drift.png
                                              Reply  Fri Apr 13 17:56:26 2018, gautam, Update, ALS, Fibers switched out 
                                                 Reply  Sat Apr 14 14:42:09 2018, gautam, Update, ALS, Fibers switched out PSLdrift.png
                                                    Reply  Tue Apr 17 14:08:29 2018, gautam, Update, ALS, Fibers switched out PSL_fluctuations.png
                                                       Reply  Fri Apr 20 00:26:34 2018, gautam, Update, ALS, Fibers switched out 
                                                          Reply  Sat Apr 21 20:25:12 2018, gautam, Update, ALS, PSL fiber pickoff status PSL-beatMouthPickoff.png
                                  Reply  Thu Apr 19 00:03:51 2018, gautam, Update, IOO, More IMC NBing IMC_NB_20180409.pdfIMC_coils_20180418.pdf
                                     Reply  Thu Apr 19 17:15:35 2018, gautam, Update, IOO, More IMC NBing IMC_NB_20180419.pdfIMC_coilDriverNoises_20180419.pdf
                                        Reply  Sun May 20 23:43:50 2018, gautam, Update, IOO, Coil driver noise re-measurement MC1_coilDriver.pdfMC2_coilDriver.pdfMC3_coilDriver.pdfMC_coilDriverNoises.tgzActuationChainNoiseMeas.pdf
                                           Reply  Tue May 22 17:26:25 2018, gautam, Update, IOO, MC1 Coil Driver pulled out 
                                              Reply  Tue May 22 23:28:01 2018, gautam, Update, IOO, MC1 Coil Driver pulled out 
                                                 Reply  Wed May 23 17:58:48 2018, gautam, Update, IOO, MC1 Coil Driver pulled out MC1_monitorTF.pngMC1_ULnoise.pdf
                                                    Reply  Wed May 30 10:17:46 2018, gautam, Update, IOO, MC1 Coil Driver pulled out EM_pickup.pdfcoilDriverNoiseComparison.pdf
                            Reply  Fri Apr 6 22:23:53 2018, Kevin, Update, IOO, Coil driver noise MC2_coil_driver.pdf
Message ID: 13770     Entry time: Thu Apr 19 17:15:35 2018     In reply to: 13765     Reply to this: 13870
Author: gautam 
Type: Update 
Category: IOO 
Subject: More IMC NBing 

Summary:

Today, I repeated the coil driver noise measurement. Now, the coil driver noise curve in the noise budget plot (Attachment #1) is the actual measurement of all 12 coils (made with G=100 SR560). I am also attaching the raw voltage noise measurement (input terminated in 50ohms, Attachment #2). Note that POX11 spectrum has now been re-measured with analog de-whitening engaged on all 3 optics such that the DAC noise contribution should be negligible compared to coil driver noise in this configuration. The rows in Attachment #2 correspond to 800 Hz span (top) and full span (bottom) on the FFT analyzer.

Details:

The main difference between this measurement, and the one I did middle of last year (which agreed with the expectation from LISO modeling quite well) is that this measurement was done in-situ inside the eurocrate box while last year, I did everything on the electronics benches. So I claim that the whole mess of harmonics seen in the raw measurements are because of some electronics pickup near 1X6. But even disregarding the peaky features, the floor of ~30nV/rtHz is ~6x than what one would expect from LISO modeling (~5nV/rtHz). I confirmed by looking that the series resistance for all 3 MC optics is 430ohms. I also did the measurement with the nominal bias voltages applied to the four channels (these come in via the slow ADCs). But these paths are low-passed by an 8th order low pass with corner @ 1Hz, so at 100 Hz, even 1uV/rtHz should be totally insignificant. I suppose I could measure (with EPICS sine waves) this low-pass filtering, but it's hard to imagine this being the problem. At the very least, I think we should get rid of the x3 gain on the MC2 coil driver de-whitening board (and also on MC1 and MC3 if they also have the x3 factor).

Attachment 1: IMC_NB_20180419.pdf  429 kB  Uploaded Thu Apr 19 18:51:09 2018  | Hide | Hide all
IMC_NB_20180419.pdf
Attachment 2: IMC_coilDriverNoises_20180419.pdf  308 kB  Uploaded Thu Apr 19 18:56:09 2018  | Hide | Hide all
IMC_coilDriverNoises_20180419.pdf
ELOG V3.1.3-